# The gauge symmetry - unbroken and broken - in the QCD ground state

### v2

@@ Abstract
The main purpose of this work is to report on convincing progress made in the investigation of the one of the long-standing conceptual problems in QCD, namely whether the gauge symmetries of the QCD Lagrangian and its ground state coincide or not? For this purpose, we use only the tensor algebra derivation rigorous rules in dealing with the corresponding skeleton loop integrals, contributing into the full gluon self-energy. It has been exactly (i.e., without making any kind of the simplifications) proven that in the QCD ground state may exist the two principally different types of the gluon field configurations. One of them maintains the QCD Lagrangian’s gauge symmetry, but another one does not.The dynamical source of this effect has to be identified with the constant tadpole term, having the dimension of a mass squared. The exact proof has been given that it cannot be disregarded from the theory and its ground state by any means. The gauge symmetry of the QCD Lagrangian forbids such term, but the perturbative renormalizability of the theory will not be affected. This allows to present the new significant insights into the gauge and dynamical structures of the true QCD ground state, and sheds light on a mass squared dynamical generation in it.
@@
theoretical particle physics, Quantum Chromodynamics, ground state, gluon full self-energy, dynamical and gauge structures.
@@ 1 Introduction
The quark model (QM) treats the strongly-interacting hadrons (baryons and mesons) as bound-states of quarks, by emitting and absorbing gluons. The theory, which purpose is to describe the properties of the observed hadrons in terms of the non-observable quarks and gluons from first principles, is Quantum Chromodynamics (QCD) 1 ; 2 ; 3 ; 4 ; 5 ; 6 ; 7 ; 8 ; 9 . It is widely accepted as the quantum gauge field theory of strong interactions. Being such a theory, it has to undergo the two phase transitions: the first one at the fundamental (microscopic) quark-gluon level is – the confinement phase transition – in order to explain why all the physical states are colour-singlets. The second one at the (macroscopic) hadronic level is – the PCAC (partially conserved axial currents) phase transition – in order to explain the soft pion physics 2 ; 23 .
@@
B. The dynamical generation of a mass squared in the vacuum of QCD, since the gauge invariance of its Lagrangian forbids such kind of terms, like the gluon mass term $M^{2}_{g}A_{\mu}A_{\mu}$.
The solutions of these fundamental problems is a key to the explanation of the first confinement phase transition in QCD. The properties and symmetries of the QCD Lagrangian, and thus including its Yang – Mills (YM) part, are well-known 1 ; 2 ; 3 ; 4 ; 5 ; 6 ; 7 ; 8 ; 9 ; 10 . The propagation of gluons is one of the main dynamical effects in the QCD ground state. The importance of the corresponding equation of motion is due to the fact that its investigation is supposed to reflect the dynamical and gauge structures of the QCD ground state. The gluon Schwinger – Dyson (SD) equation 2 ; 10 ; 11 ; 12 ; 13 ; 14 ; 15 ; 16 ; 17 ; 18 ; 19 ; 20 is highly non-linear (NL) because of the self-interaction of massless gluon modes. They are taken into account by the full gluon self-energy, which is the main part of the gluon SD equation itself. We will show that the general properties of the full gluon self-energy point out on some new dynamical and gauge aspects of the true QCD ground state. Due to these inputs, the novel insights into a mass dynamical generation in the QCD vacuum are also present within proposed below formalism, based on the tensor algebra derivation rigorous rules only.
2 The gluon SD equation
The structure of the gluon SD equation is given in this section in some necessary details. Analytically it looks like
@@ 2 The gluon SD equation
Figure 1: The SD equation for the full gluon propagator as present in 10 .
@@
where $\Pi^{q}_{\rho\sigma}(q)$ describes the skeleton loop contribution of the quark degrees of freedom as an analogue to the vacuum polarization tensor in Quantum Electrodynamics (QED) 7 ; 21 . Note that here and below the superscript or subscript ’$q$’ means quark (not to be mixed up with the gluon momentum variable $q$). The gluon part of the full gluon self-energy by itself is the sum of a few independent terms as follows:
$\Pi^{g}_{\rho\sigma}(q;D)=\Pi^{gh}_{\rho\sigma}(q)+\Pi^{(1)}_{\rho\sigma}(q;D^{2})+\Pi^{(2)}_{\rho\sigma}(q;D^{4})+\Pi^{(2^{\prime})}_{\rho\sigma}(q;D^{3}),$ (2.3)
and $\Pi^{gh}_{\rho\sigma}(q)$ describes the skeleton loop contribution associated with the ghost degrees of freedom. $\Pi^{(1)}_{\rho\sigma}(q;D^{2})$ represents the skeleton loop contribution, containing the 3-gluon vertices only. Finally, $\Pi^{(2)}_{\rho\sigma}(q;D^{4})$ and $\Pi^{(2^{\prime})}_{\rho\sigma}(q;D^{3})$ describe the skeleton two-loop contributions, which combine the 3- and 4-gluon vertices. All these quantities are given by the corresponding skeleton loop diagrams in Fig. 1, and they are independent from each other. The last four terms explicitly contain the full gluon propagators in the corresponding powers symbolically shown above. They form the NL part of the gluon SD equation. The analytical expressions for the corresponding skeleton loop integrals 22 , in which the symmetry and combinatorial coefficients and signs have been included, are not important here. We are not going to calculate any of them explicitly, and thus to introduce into them any truncations/approximations/assumptions or choose some special gauge. Any skeleton loop integral in Fig. 1 is the sum of the infinite number of terms. Moreover, the full vertices entering these skeleton loop integrals are themselves determined by an infinite series of the corresponding multi-loop skeleton terms 2 ; 11 ; 12 ; 13 ; 14 ; 15 ; 16 ; 17 ; 18 ; 19 ; 20 . In these NL series the dependence on the coupling constant may not be simple, i.e., in fact, it is not known. Such kind of series are the so-called ’cluster’ expansions 23 , indeed.
The constant tadpole term $\Pi_{\rho\sigma}^{t}(D)$ is defined as follows:
$\Pi_{\rho\sigma}^{t}(D)\sim\int d^{4}lD_{\alpha\beta}(l)T^{0}_{\rho\sigma\alpha\beta}=g_{\rho\sigma}\Delta^{2}_{t}(D)=[T_{\rho\sigma}(q)+L_{\rho\sigma}(q)]\Delta^{2}_{t}(D),$ (2.4)
where $L_{\rho\sigma}(q)=q_{\rho}q_{\sigma}/q^{2}$, so that the tadpole term contributes into the both transverse and longitudinal components of the full gluon propagator (2.1). In what follows the constant $\Delta^{2}_{t}(D)$ will be called as the tadpole term itself, since it deviates from the definition (2.4) by the tensor $g_{\rho\sigma}$ only.
@@
$D_{\mu\nu}(q)=D^{0}_{\mu\nu}(q)$ $+$ $D^{0}_{\mu\rho}(q)i\left[\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D)\right]D_{\sigma\nu}(q)$ (2.5)
$+$ $D^{0}_{\mu\rho}(q)ig_{\rho\sigma}\Delta^{2}_{t}(D)D_{\sigma\nu}(q).$
@@
All the terms which contribute to the full gluon self-energy eq. (2.2), and hence eq. (2.5), are tensors, having the dimensions of a mass squared. All these skeleton loop integrals are therefore quadratically divergent (QD), i.e., ultraviolet (UV) divergent in the perturbation theory (PT) regime, and so they are assumed to be regularized. Contrary to QED, QCD being a non-abelian gauge theory can suffer from the severe infrared (IR) singularities in the $q^{2}\rightarrow 0$ limit (i.e., more singular than the PT one, which is proportional to $\sim 1/q^{2}$). Evidently, they will be generated due to the self-interaction of massless gluon modes. Thus, all the possible subtractions at zero may be dangerous 2 ; 3 ; 4 ; 5 ; 6 ; 7 . That is why in all the quantities below, the dependence on the finite (slightly different from zero) dimensionless subtraction point $\alpha$ has to be understood. In other words, all the subtractions at zero and the Taylor expansions around zero should be understood as the subtractions at $\alpha$ and the structure of the Taylor expansions near $\alpha$, where they are justified to be used. From the technical point of view, however, it is convenient to put formally $\alpha=0$ in all the expressions and derivations below, and to restore the explicit dependence on non-zero $\alpha$ in all the quantities only at the later stage. In all the quantities where the dependence on the dimensionless UV regulating parameter $\tilde{\lambda}$ and $\alpha$ is not shown explicitly, nevertheless, it should be also assumed. For example, $\Pi_{\rho\sigma}(q;D)\equiv\Pi_{\rho\sigma}(q;D,\tilde{\lambda},\alpha)$ and similarly for all other quantities. This means that all the expressions are regularized (i.e., we render infinite skeleton loop integrals to be finite).
@@
The tadpole term $\Delta^{2}_{t}(D)$, defined in eq. (2.4), is nothing else but the QD skeleton loop integral, which is already regularized from below and above as well as all the other such kind of constants which may appear in the theory, as a results of the proper subtraction scheme (see below). Being a mass squared regularized quantity, it is explicitly present in the QCD ground state, see eq. (2.5). However, it violets the gauge symmetry of the QCD Lagrangian. Then a natural question arises why is it present in the vacuum of QCD at all if it makes theory to look un- renormalizable from the very beginning? The standard solution to this dilemma was to remove this and all other QD constants in order to make theory renormalizable in accordance with the QCD Lagrangian’s exact gauge symmetry. However, here we are going to solve this dilemma by asking the question is it possible to retain the tadpole term in the QCD vacuum, but without affecting the PT renormalization properties of the theory? This is important to ask because, unlike to other QD constants, the tadpole term is only one which generates a mass squared scale parameter 6 . Also, one of our primary aims will be to clarify its role in the dynamical and gauge structures of the QCD vacuum in more detail. For this purpose we will develop a suitable formalism, based only on the rigorous rules of the tensor algebra derivations.
@@
3 Slavnov-Taylor identities
The basic relation to which the full gluon propagator should satisfy is the corresponding Slavnov-Taylor (ST) identity 2 ; 24 ; 25 , namely
@@ 3 Slavnov-Taylor identities
where the invariant function $d(q^{2})$ is the corresponding Lorentz structure of the full gluon propagator (or the gluon invariant function). Throughout this paper we use the standard definition of $T_{\mu\nu}(q)=\delta_{\mu\nu}-q_{\mu}q_{\nu}/q^{2}=\delta_{\mu\nu}-L_{\mu\nu}(q)$ in Euclidean metric. Any invariant functions associated with the projective operators $T_{\mu\nu}(q)$ and $\delta_{\mu\nu}$ are the same, and thus $D_{\mu\nu}(q)$ is defined up to its longitudinal part $L_{\mu\nu}(q)$. This should be also valid for its equation of motion.
If one neglects all the contributions to the full gluon self-energy in eq. (2.1), by setting $d(q^{2})=1$ in eq. (3.2), then one obtains the free gluon propagator
@@
Let us note in advance that from the very beginning the gauge-fixing parameter for the full gluon propagator $\xi$ is not to be identified with its free counterpart $\xi_{0}$. Under which conditions they are the same or different is one of the main subjects of our investigation in this paper. Since the gauge freedom in QCD is determined by $\xi_{0}$, the gauge-fixing parameter $\xi$ has to be a function of $\xi_{0}$ i.e., $\xi=f(q^{2};\xi_{0})$ and called the gauge-fixing function.
Emphasizing, these and all other ST identities which appear in QCD 2 ; 3 ; 4 ; 5 ; 6 ; 7 ; 9 ; 11 ; 12 ; 13 ; 14 ; 15 ; 16 ; 17 ; 18 ; 19 ; 20 are important for the general renormalizability of the theory, and thus they ”are exact constraints on any solution to QCD” 2 .
3.1 Transverse relations for the full gluon self-energy
@@
Let us now perform some algebraic derivations which are necessary for the further purposes. Contracting the full gluon SD eq. (2.1) and eq. (2.5) with $q_{\mu}$ and $q_{\nu}$, on account of the relations (3.1)-(3.4), one gets
$q_{\rho}q_{\sigma}\Pi_{\rho\sigma}(q;D)$ $=$ ${(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2},$
$q_{\rho}q_{\sigma}[\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D)]$ $=$ ${(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2}-q^{2}\Delta^{2}_{t}(D),$ (3.5)
@@
and they related to each other through the relation (2.2) as well as eq. (2.1) is related to eq. (2.5), which is as it should be. The first line in the transverse relations (3.5) is the tensor relation, while its second one is nothing else but the tensor identity. The second transverse relation explicitly depends on the tadpole term $\Delta^{2}_{t}(D)$. If it is formally put zero, then the both transverse relations (3.5) coincide with each other. This underlines its important role in the QCD ground state, which one precisely is to be clarified in our paper.

### v3

# The gauge symmetry violation in the QCD ground state

@@ Abstract
The purpose of this work is to report on convincing progress made in the investigation of the true dynamical and gauge structures of the QCD ground state (vacuum). It also sheds new significant insights into a mass squared dynamical generation in it. For this purpose, we use only the tensor algebra derivation rigorous rules in dealing with the corresponding skeleton loop integrals, contributing into the full gluon self-energy. The novel constraint on any solution to QCD has been exactly and uniquely derived in the framework of the the Slavnov-Taylor identities for the gauge particles propagators. Its general solution makes it possible to prove that the gauge symmetry of the QCD Lagrangian is broken in its ground state. The dynamical source of this effect has to be identified with the constant tadpole term, having the dimension of a mass squared. The exact proof has been given that it cannot be disregarded from the theory and its ground state by any means. The gauge symmetry of the QCD Lagrangian forbids such kind of terms, but the perturbative renormalizability of the theory will not be affected. The particular solution of the novel constraint describes the gluon field configurations with unbroken gauge symmetry in the QCD vacuum. The solution with broken gauge symmetry dominates over the solution with unbroken one at large distances, while being in agreement with it at short distances.
@@
Theoretical particle physics, Quantum Chromodynamics, ground state, full gluon self-energy, gauge
@@ 1 Introduction
The quark model (QM) treats the strongly-interacting hadrons (baryons and mesons) as bound-states of quarks, by emitting and absorbing gluons. The theory, which purpose is to describe the properties of the observed hadrons in terms of the non-observable quarks and gluons from first principles, is Quantum Chromodynamics (QCD) [1, 2, 3, 4, 5, 6, 7, 8, 9]. It is widely accepted as the quantum gauge field theory of strong interactions. Being such a theory, it has to undergo the two phase transitions: the first one at the fundamental (microscopic) quark-gluon level is – the confinement phase transition – in order to explain why all the physical states are colour-singlets. The second one at the (macroscopic) hadronic level is – the PCAC (Partially Conserved Axial Currents) phase transition – in order to explain the soft pion physics [2, 23].
@@
B. The dynamical generation of a mass squared in the vacuum of QCD, since the gauge invariance of its Lagrangian forbids such kind of terms, for example like the gluon mass term $M^{2}_{g}A_{\mu}A_{\mu}$.
The solutions of these fundamental problems is a key to the explanation of the first confinement phase transition in QCD. The properties and symmetries of the QCD Lagrangian, and thus including its Yang – Mills (YM) part, are well-known [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. The propagation of gluons is one of the main dynamical effects in the QCD ground state. The importance of the corresponding equation of motion is due to the fact that its investigation is supposed to reflect the dynamical and gauge structures of the QCD ground state. The gluon Schwinger – Dyson (SD) equation [2, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20] is highly non-linear (NL) because of the self-interaction of massless gluon modes. They are taken into account by the full gluon self-energy, which is the main part of the gluon SD equation itself. We will show that the general properties of the full gluon self-energy point out on some new dynamical and gauge aspects of the true QCD ground state. Due to these impacts, the novel insights into a mass dynamical generation in the QCD vacuum are also present within proposed below formalism. It is based on the tensor algebra derivation rigorous rules only, which is widely using in the theoretical and mathematical physics.
2 Gluon SD equation
The system of the SD equations of motions in QCD, including the above-mentioned gluon SD equation, in general, contains much more informations on the properties of the theory than its Lagrangian can provide at all [2]. So it might be the place where a mas squared scale parameter, forbidden by the QCD Lagrangian gauge symmetry, could be generated. Among them, the special role belongs to the equations, describing the propagation of the fundamental constituents in the QCD vacuum – gluons and quarks. The gluon SD equation has a rather complicated tensor structure because of its above-mentioned NL character. Let us emphasize that all the known interactions in nature (electroweak, gravitational and strong) are described by the gauge theories. The common mathematical language for them is the tensor algebra, which is only one used throughout this paper. That is why our works will be easily understood by a broad spectrum of researches from the different areas of the theoretical physics. The gluon SD equation analytically looks like
@@ 2 Gluon SD equation
Figure 1: The SDE for the full gluon propagator, as present in [10].
@@
where $\Pi^{q}_{\rho\sigma}(q)$ describes the skeleton loop contribution of the quark degrees of freedom as an analogue to the vacuum polarization tensor in Quantum Electrodynamics (QED) [7, 21]. Here and below the superscript or subscript ’$q$’ means quark (not to be mixed up with the gluon momentum variable $q$). The gluon part of the full gluon self-energy by itself is the sum of a few independent terms as follows:
$\Pi^{g}_{\rho\sigma}(q;D)$ $=$ $\Pi^{gh}_{\rho\sigma}(q)+\Pi^{(1)}_{\rho\sigma}(q;D^{2})$ (2.3)
$+$ $\Pi^{(2)}_{\rho\sigma}(q;D^{4})+\Pi^{(2^{\prime})}_{\rho\sigma}(q;D^{3}),$
and $\Pi^{gh}_{\rho\sigma}(q)$ describes the skeleton loop contribution associated with the ghost degrees of freedom. $\Pi^{(1)}_{\rho\sigma}(q;D^{2})$ represents the skeleton loop contribution, containing the 3-gluon vertices only. Finally, $\Pi^{(2)}_{\rho\sigma}(q;D^{4})$ and $\Pi^{(2^{\prime})}_{\rho\sigma}(q;D^{3})$ describe the skeleton two-loop contributions, which combine the 3- and 4-gluon vertices. All these quantities are given by the corresponding skeleton loop diagrams in Fig. 1, and they are independent from each other. The last four terms explicitly contain the full gluon propagators in the corresponding powers symbolically shown above. The analytical expressions for the corresponding skeleton loop integrals [22], in which the symmetry and combinatorial coefficients and signs have been included, are not important here. We are not going to calculate any of them explicitly, and thus to introduce into them any truncations/approximations/assumptions or choose some special gauge. Any skeleton loop integral in Fig. 1 is the sum of the infinite number of terms. Moreover, the full vertices entering these skeleton loop integrals are themselves determined by an infinite series of the corresponding multi-loop skeleton terms [2, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. In these NL series the dependence on the coupling constant may not be simple, i.e., in fact, it is not known. Such kind of series are the so-called ’cluster’ expansions [23], indeed.
The constant tadpole term $\Pi_{\rho\sigma}^{t}(D)$ is defined as,
$\Pi_{\rho\sigma}^{t}(D)\sim\int d^{4}lD_{\alpha\beta}(l)T^{0}_{\rho\sigma\alpha\beta}=g_{\rho\sigma}\Delta^{2}_{t}(D),$ (2.4)
where $g_{\rho\sigma}=T_{\rho\sigma}(q)+L_{\rho\sigma}(q)=T_{\rho\sigma}(q)+q_{\rho}q_{\sigma}/q^{2}$. In what follows the constant $\Delta^{2}_{t}(D)$ will be called as the tadpole term itself, since it deviates from the definition (2.4) by the tensor $g_{\rho\sigma}$ only.
@@
$D_{\mu\nu}(q)=D^{0}_{\mu\nu}(q)+D^{T}_{\mu\rho}(q)+D^{L}_{\mu\rho}(q)$ (2.5)
with
$D^{T}_{\mu\nu}(q)$ $=$ $D^{0}_{\mu\rho}(q)i\left[\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D)\right]D_{\sigma\nu}(q),$
$D^{L}_{\mu\nu}(q)$ $=$ $D^{0}_{\mu\rho}(q)ig_{\rho\sigma}\Delta^{2}_{t}(D)D_{\sigma\nu}(q).$ (2.6)
@@
All the terms which contribute to the full gluon self-energy eq. (2.2), and hence eq. (2.5), are tensors, having the dimensions of a mass squared. All these skeleton loop integrals are therefore quadratically divergent (QD), i.e., ultraviolet (UV) divergent in the perturbation theory (PT) regime, and so they are assumed to be regularized from above. Contrary to QED, QCD being a non-abelian gauge theory can suffer from the severe infrared (IR) singularities in the $q^{2}\rightarrow 0$ limit (i.e., more singular than the PT one, which is proportional to $\sim 1/q^{2}$). Evidently, they will be generated due to the self-interaction of massless gluon modes. Thus, all the skeleton loop integrals, contributing to the full gluon self-energy, are assumed to be regularized from below as well. This means that all the expressions are regularized (i.e., we render infinite skeleton loop integrals to be finite). How precisely the regulating parameters have to be introduced is not important here, since we are not going to calculate any of these skeleton loop integrals explicitly, as emphasized above. They will be assumed but not shown in all regularized expressions, for simplicity.
@@
The tadpole term $\Delta^{2}_{t}(D)$, defined in eq. (2.4), is nothing else but the QD skeleton loop integral, which is already regularized from below and above as well as all the other such kind of constants which may appear in the theory by any possible ways (the subtraction schemes, the iteration solutions, etc., see below). Being a mass squared regularized quantity, it is explicitly present in the QCD ground state, see eq. (2.5). However, it violets the gauge symmetry of the QCD Lagrangian. Then a natural question arises why is it present in the vacuum of QCD at all if it makes the theory to look non-renormalizable from the very beginning? The standard solution to this dilemma was to remove this and all other QD constants in order to make theory renormalizable in accordance with the QCD Lagrangian’s exact gauge symmetry. However, here we are going to solve this dilemma by asking the question is it possible to retain the tadpole term in the QCD vacuum, but without affecting the PT renormalization properties of the theory? This is important to ask because, unlike to other QD constants, the tadpole term is only one which dynamically generates a mass squared scale parameter, dominated by the point-like four-gluon vertex only, and thus does not depending on the external gluon momentum [6]. Also, one of our primary aims will be to clarify its role in the dynamical and gauge structures of the QCD vacuum in more detail.
@@
3 Exact constraint on any solution to QCD
It is well known that all the Slavnov-Taylor (ST) identities which appear in QCD [1] are important for the renormalizability of the theory, and thus they ”are exact constraints on any solution to QCD” [2]. Among them the ST identities [2, 24, 35] for the gauge particle propagators play a dominant role, since they determine the gauge structure of the QCD ground state (vacuum).
The ST identity for the full gluon propagator looks like
@@ 3 Exact constraint on any solution to QCD
where the invariant function $d(q^{2})$ is the corresponding Lorentz structure of the full gluon propagator (or the gluon invariant function). Throughout this paper we use the standard definition of $T_{\mu\nu}(q)=\delta_{\mu\nu}-q_{\mu}q_{\nu}/q^{2}=\delta_{\mu\nu}-L_{\mu\nu}(q)$ in Euclidean metric. Any invariant functions associated with the projection operators $T_{\mu\nu}(q)$ and $\delta_{\mu\nu}$ are the same, and thus $D_{\mu\nu}(q)$ is defined up to its longitudinal part $L_{\mu\nu}(q)$. This should be also valid for its equation of motion. By formally setting $d(q^{2})=1$ in eq. (3.8), then one obtains the free gluon propagator
@@
Let us note in advance that from the very beginning the gauge-fixing parameter for the full gluon propagator $\xi$ is not to be identified with its free counterpart $\xi_{0}$. Since the gauge freedom in QCD is determined by $\xi_{0}$, there should exist a relation between them.
@@
Let us now perform some exact algebraic derivations which are necessary for the further purposes. Contracting the full gluon SD eq. (2.1) with $q_{\mu}$ and $q_{\nu}$, on account of the relations (3.7)-(3.10), one gets
$q_{\rho}q_{\sigma}\Pi_{\rho\sigma}(q;D)={(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2},$ (3.11)
where the full gluon self-energy, due to the eqs. (2.2) and (2.4), can be present as follows:
$\Pi_{\rho\sigma}(q;D)=\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D)+\delta_{\rho\sigma}\Delta^{2}_{t}(D),$ (3.12)
and contracting it with with $q_{\rho}$ and $q_{\sigma}$, and because of the previous relation (3.11), one obtains
$q_{\rho}q_{\sigma}[\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D)]={(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2}-q^{2}\Delta^{2}_{t}(D).$ (3.13)
@@
Obviously, this transverse relation can be obtained by the corresponding contraction of eq. (2.5). It explicitly depends on the tadpole term $\Delta^{2}_{t}(D)$. If it is formally put zero, i.e., by hand, then the both transverse relations (3.11) and (3.13) coincide with each other. This underlines its important role in the QCD ground state, which one precisely is to be clarified in our paper. The transverse relation (3.11) and the transverse relation obtained from eq. (3.12) are independent from each other, since the first one has been derived by using the gluon SD eq. (2.2), while the second one by using the initial definition (2.6). From all these transverse relations one concludes that by themselves they cannot remove the QD constants which may appear in the theory (in order to make it in the PT sense renormalizable), and to fix the relation between $\xi$ and $\xi_{0}$.
Let us now remind that in QCD the quark skeleton loop contribution into the full gluon self-energy can be always made transverse because of the electric charge conservation which flows around the closed loop (in complete analogy with QED, see [10] and derivations below). Therefore, $q_{\rho}q_{\sigma}\Pi^{q}_{\rho\sigma}=0$ and this relation holds independently from the gluon contribution, which has been defined in eq. (2.3). At the same time, it is well-known that in QCD that just the Faddeev – Popov (FP) ghost contribution [43], $\Pi^{gh}_{\rho\sigma}(q)$ makes the transverse relation for this term satisfied (independently from the tadpole term (2.4), which itself has the transverse projection operator structure). It is instructive to emphasize that each term in the sum (2.3) cannot be made transverse, only the whole sum. How explicitly this works can be found in any text book on QCD. For the most transparent calculations up to one loop contribution to the transverse gluon self-energy (2.3) see for example [3, 4, 5, 6]. Of course, such a transversality will be achieved in the higher loop iteration terms as well, i.,e, it is a general fact, as it is requested by the satisfied transverse relation for the (2.3) term. It is a sum of the skeleton loop contributions, as underlined from the very beginning in this work. This means that $q_{\rho}q_{\sigma}\Pi^{g}_{\rho\sigma}(q;D)=0$ and this relation holds independently from the quark contribution.
So that, in general, in QCD one has
$q_{\rho}q_{\sigma}\Pi^{q}_{\rho\sigma}(q)=q_{\rho}q_{\sigma}\Pi^{g}_{\rho\sigma}(q;D)=0,$ (3.14)
and then from the relation (3.13), one immediately arrives at the exact condition, namely
${(\xi_{0}-\xi)\over\xi\xi_{0}}q^{2}=\Delta^{2}_{t}(D),$ (3.15)
which can be treated as the exact constraint on any solution to QCD, since coming out from the corresponding ST identities, and deriving before going to the formulation of any proper subtraction scheme.This underlines its intrinsic, independent and important status in the theory.
3.1 General solution
The general solution of the exact constraint (3.15) determines the function $\xi=f(q^{2};\xi_{0})$ as follows:
$\xi=f(q^{2};\xi_{0})={\xi_{0}q^{2}\over q^{2}+\xi_{0}\Delta^{2}_{t}(D)}.$ (3.16)
We will call it as the generalized gauge-fixing function approach to QCD as well. It has been exactly (i.e., without making any kind of the simplifications) and uniquely defined in a such new manner for the first time.This expression demonstrates the NL dependence $\xi=f(q^{2};\xi_{0})$ on $\xi_{0}$. The linear relation $\xi\rightarrow\xi_{0}$ will be recovered in the PT $q^{2}\rightarrow\infty$ limit, when the ratio $(\Delta^{2}_{t}(D)/q^{2})$ is to be suppressed in this regime at finite $\xi_{0}$. We have established the general boundary condition in QCD for the behaviour of the gauge-fixing function in the PT limit for the regularized full gluon propagator. Since the general solution (3.16) has been obtained within the ST identities formalism, we will call it as the general solution to QCD. Let us underline that to know it means to know the true gauge structure of the QCD vacuum and its theory at the fundamental quark-gluon level (for more detailed description see section 5 as well).
3.2 Particular solution
The exact constraint (3.15) has also a particular solution, namely